Current position: S72 13′ 30.5″, E023 25′ 00.2″
Today’s action: Met obs. Snow sampling. Eight fuel scoots in two loads of four were taken down the descent into the valley which splits the mountains. The lower part of this route requires a braking Cat at the rear of each load. No problem with the route and it is good to get the tracks laid whilst the weather is so good.
Intentions tomorrow: No further load movements tomorrow. We will explore the local area on foot/ski if conditions are suitable. Cold finger tests (White Mars).
Long term intentions: We will keep camp in this current position for several more days. Weather permitting we will explore the local area further on foot/ski.
Comments are closed